f(x)=(x^2-2x+1)(x^3-1) find all the x-values where the slope of the tangent line is 3
Answer #1

(Date: IL tal= (2-2x+1) (231) =_ (23-1) - 24173-) + 173-1) f(W) = xs-x² - 204 + 2x + 23-) f(x) = x 214 + x - x2 +2n-1 to resp